### v1

# Boosting Graph Foundation Model from Structural Perspective

Abstract
Graph foundation models have recently attracted significant attention due to its strong generalizability. Although existing methods resort to language models to learn unified semantic representations across domains, they disregard the unique structural characteristics of graphs from different domains. To address the problem, in this paper, we boost graph foundation model from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the super nodes fuse the information of anchor nodes and class labels, where each anchor node captures the information of a node or a graph instance to be classified. Instead of using the raw graph structure, we connect super nodes to all nodes within their neighborhood by virtual edges. This new structure allows for effective information aggregation while unifying cross-domain structural characteristics. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Experimental results on various datasets and tasks demonstrate the superior performance of BooG. We provide our code and data here: https://anonymous.4open.science/r/BooG-EE42/.
1 Introduction
Graph learning has achieved outstanding results in a wide range of application fields in recent years, including social networks [22, 13], citation networks [18], molecular graphs [2, 33] and recommendation systems [6]. For graph data containing complex structural information, graph neural networks (GNNs) [25, 11, 29] leverage designed message-passing mechanisms to learn consistent representations for similar instances and apply them to various downstream tasks, such as node classification [11], graph classification [29] and link prediction [31].
Although general GNN models have achieved widespread success, they often face limitations due to training labels and can only perform a single task. In practical applications, accurate ground truth labels often come at a high cost [9]. To address these challenges, graph self-supervised learning (SSL) [26, 30, 9] utilizes the structure or characteristics of the data itself for learning, without the need for explicitly providing labels for the training instances. For example, graph contrastive learning (GCL) [30] supplements supervised signals by comparing the similarity and dissimilarity between different instances in the graph data. Graph SSL methods typically employ a learning pipeline involving “pre-training” and “fine-tuning”. By designing appropriate pre-training tasks, the model is enabled to generalize to downstream tasks after fine-tuning.
Although graph SSL has been demonstrated to effectively learn representations of graph data during the pre-training phase, the disparity between pre-training tasks and downstream tasks can result in suboptimal performance [16]. With the development of large language models (LLMs), prompt learning has been proposed to narrow the gap between pre-training and downstream tasks. Recently, “prompt” has been introduced into graph learning [3, 17, 23]. Such efforts aim to narrow the gap between pre-training and downstream tasks by designing unified pre-training and prompting templates. However, graph prompt learning fails to assist pre-trained models in improving domain generalization capabilities. This is because graph data from different domains often contain specific semantic characteristics. For instance, features of molecular graphs are typically vectors representing indices of nominal features of atoms, while features of citation networks often consist of bag-of-words vectors describing paper titles and abstracts.
To enhance the generalization capability of graph learning for cross-domain graph tasks, graph foundation models based on LLMs aim to learn unified semantic representations on graph data from different domains by text-attribute graphs (TAGs) [14, 28, 7]. However, these approaches are still constrained by the unique structural characteristics of graph data from different domains. For example, citation networks contain citation relationships between academic papers, often exhibiting high levels of clustering, where papers referencing each other may involve similar topics or domains. On the other hand, chemical molecular graphs typically exhibit fixed topological structures and physical-chemical properties. The structural characteristics from different domains make it challenging for models to generalize the learned structural knowledge to other domains. To further enhance the model generalizability, there naturally arises a question: Can we develop a graph foundation model that unifies cross-domain structural characteristics?
To address the aforementioned challenges, in this paper, we propose the Boosting Graph Foundation Model, namely, BooG, which achieves robust performance and generalization capabilities across various graph domains and tasks. First, BooG uses TAGs to map textual descriptions of input data to a unified space by a pre-trained language model (LM). It then introduces virtual super nodes to construct unified structural characteristics for graph data from different domains, addressing the challenge of domain-specific structural characteristics that hinders generalization. Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified. Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges. Details will be illustrated in Figure 3 later. This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains. In this way, cross-domain model generalization can be boosted from structural perspective. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. We conduct extensive experiments on six datasets with diverse domains and tasks. In most cases, our proposed approach leads the state-of-the-art competitors. We also conduct ablation studies to validate the necessity of each component in BooG.
2 Related Works
Graph Neural Network.
As an effective technique for handling graph-structured data, GNNs [25, 11, 29, 12, 4, 27] can effectively capture complex patterns based on input node features and structural information. For instance, graph attention networks (GAT) [25] utilize attention mechanisms to learn the importance of neighboring nodes and aggregate their information with learned weights. Graph Isomorphism Network (GIN) [29] adopts a unique graph encoding method that emphasizes distinguishing non-isomorphic structures, enhancing the model’s representational capacity. However, the effectiveness of GNNs often relies on a large number of training labels, implying high labeling costs. Additionally, the design of GNNs is often tailored to specific tasks, making it challenging to generalize to other downstream tasks. For instance, GNNs designed for node classification tasks often focus more on local information of the instances, while methods designed for graph classification tasks require more attention to the global information of the instances.
Graph Self-Supervised Learning.
Graph SSL [26, 30, 9] is primarily used to address the problem of lacking labels in graph learning tasks. These methods typically employ learning pipelines that include “pre-training” and “fine-tuning” stages. Through carefully designed contrastive learning or generative pre-training tasks, the model can generalize the learned knowledge to downstream tasks after fine-tuning. For example, GraphCL [30] utilizes the pre-training of graph models by applying a self-supervised contrastive learning task on learned node embeddings. DGI [26] introduces a self-supervised pre-training task aimed at maximizing the mutual information between local node views and global graph views. While graph SSL has been demonstrated to effectively learn instance representations during the pre-training phase, the disparity between pre-training tasks and downstream tasks results in suboptimal performance.
Inspired by advancements in the field of NLP, graph prompting learning [3, 17, 23] has been proposed to bridge the gap between pre-training and downstream tasks. For example, GraphPrompt [17] introduces a learnable prompt layer that can automatically identify key information in the pre-trained model for downstream tasks. However, graph prompt learning fails to assist pre-trained models in improving domain generalization capabilities.
Graph Foundation Model based on LLM.
The advent of foundation models first became a reality in NLP due to the development of LLMs. Foundation models are trained on extensive data and can adapt to a wide range of data and downstream tasks [15]. Existing graph foundation models [14, 7, 28] that integrate LLMs can generally be divided into two categories: (1) Methods with LLMs as the backbone perform various graph tasks by converting graphs into text or tokens for LLMs utilization. (2) Methods with GNN as the backbone typically use pre-trained LLMs to obtain unified representations for instances, enhancing the representation capacity of GNNs and improving generalization. For example, UniGraph [7] proposes a cascade architecture of LM and GNN, along with a self-supervised training objective based on Masked Graph Modeling (MGM), and introduces graph instruction fine-tuning using LLMs to achieve zero-shot prediction capability. OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs. However, these approaches are still constrained by the structural characteristics of graph data from different domains.
3 Preliminaries
3.1 Text-Attributed Graphs
[…]
3.2 Learning scenarios.
In this paper, we propose a cross-domain and cross-task graph foundation model BooG. This model consists of two parts: a pre-trained encoder and a classifier implemented with an MLP. We first pre-train the encoder and freeze the pre-trained model’s parameters when generalizing to downstream tasks. To train the classifier on downstream tasks, we split the dataset into the train, validation, and test sets, denoted as $\mathcal{D}_{train}$, $\mathcal{D}_{val}$, and $\mathcal{D}_{test}$, respectively. Their label sets are denoted as $\mathcal{Y}_{train}$, $\mathcal{Y}_{val}$, and $\mathcal{Y}_{test}$. To evaluate the generalizability of BooG, we conduct experiments in three different problem settings and adopt three graph-related tasks including node classification, graph classification and link prediction. In supervised learning setting, the classifier will be trained on the training set $\mathcal{D}_{train}$, with the optimal parameters selected based on the validation set $\mathcal{D}_{val}$, and its performance will be finally evaluated on the test set $\mathcal{D}_{test}$. For few-shot learning, $N$-way $K$-shot tasks evaluate the in-context learning ability to apply the pre-learned knowledge to a new domain and task with $N$ classes, each represented by only $K$ labeled examples. For zero-shot learning, $K$ is set to 0, indicating no prior exposure to support instances from the target classes. This setting aims to evaluate a pre-trained model’s ability to generalize and apply its learned knowledge to unseen data categories.
Figure 1: The overall process of BooG. The model’s input includes text attribute graphs and class descriptions. BooG first utilizes a LM to unify different graph data and standardizes the input for node-level and graph-level tasks as sub-graphs. Subsequently, BooG introduces constructed super nodes into sub-graphs and learns to aggregate effective information within the sub-graphs. We freeze the parameters of the pre-trained model and provide the final instance representations through similarity matching. In particular, the similarity matching process can serve as zero-shot learning to predict unseen instances. For supervised learning and few-shot learning scenarios, BooG freezes the parameters of the pre-trained model and generalizes the capabilities of the pre-trained model to multiple downstream tasks by adjusting the parameters of the MLP.
4 Methodology
We propose a graph foundation model, BooG, which enhances the model generalizability across various graph domains and downstream tasks. In this section, we first introduce how to map the input TAGs from different domains to a unified feature space using a pre-trained language model. Next, we introduce the pre-training algorithm with the goal of learning expressive representations. Finally, we describe how to generalize the pre-trained models to multiple downstream graph tasks in different scenarios. The overall framework of BooG is summarized in Figure 1.
Figure 2: The text format for graph node and class node on Pubmed.
4.1 Unifying Graphs from Diverse Domains
Most GNNs have been designed for specific domains and relied heavily on pre-processed vector features as input, thereby posing challenges in terms of broader applicability and generalization. In this paper, inspired by previous works [14, 32], we introduce TAGs to describe graph nodes using raw texts, as illustrated in Figure 2. Even though the text description comes from different graph domains, we can apply a pre-trained LM to encode these texts into the same semantic space. And in BooG, we adopt Sentence Transformer [21], which has been shown to learn effective sentence embeddings. Let $t_{v}$ denote the text attributes of a graph node $v$, then it can be encoded by the pre-trained LM as follows:
$x(v)=LM(t_{v})\in\mathbb{R}^{d},$ (1)
where $x(v)$ is the output of the LM, and $d$ is the dimension of the output vector. Similarly, let $t_{l}$ be the text attributes of a class node $l$, then it can be encoded by the same pre-trained LM as follows:
$c(l)=LM(t_{l})\in\mathbb{R}^{d}.$ (2)
4.2 Unifying Structure from Diverse Domains
With the advancement of NLP, some works [14, 7] have introduced LMs into graph learning. By mapping graph data from different domains into a unified feature space through TAGs, these efforts aim to enhance the model’s generalization capabilities. However, graph data from different domains often contain unique structural information, posing challenges for existing methods to generalize across different domains. To address the challenges, we propose BooG, which introduces virtual super nodes to construct unified structural space for graph data from different domains.
Sub-graph. Considering that the input graph dataset could consist of a set of nodes for node classification and link prediction tasks, or a set of graphs for graph classification task, we unify the input graph data into sub-graphs composed of anchor nodes and their neighborhoods. Formally, we define a sub-graph $\mathcal{G}_{sub}=\left\{s,\mathcal{N}(s)\right\}$, where $s$ is an anchor node representing a node or a graph instance, and $\mathcal{N}(s)$ denotes its neighborhoods. For node-level tasks, the anchor node $s$ represents an arbitrary node in the input graph, and $\mathcal{N}(s)$ denotes the $k$-hop neighbor nodes of the anchor node $s$. The representation of the anchor node $s$ is the node features encoded by LM, i.e., $h(s)=x(s)$. For graph-level tasks, an anchor node $s$ represents a graph, so the representation of an anchor node $s$ is defined as $h(s)=R_{graph}(\mathcal{G})$, where $R_{graph}$ is a pooling function and the neighborhoods $\mathcal{N}(s)$ include all the nodes in the graph.
Figure 3: The super nodes in sub-graphs.
Super Node. We construct virtual super nodes by fusing each anchor node with all class labels, as shown in Figure 3. For each input node/graph instance, there is only one anchor node $s$ and $C$ super nodes $p$, where $C=\left|\mathcal{L}\right|$ represents the number of classes in a graph dataset. The representation of the $j$-th super node for the $i$-th node/graph can be computed as follows:
$h(p_{ij})=h(s_{i})+\alpha c(l_{j}),j\in\left[1,C\right],$ (3)
where $h(p_{ij})\in\mathbb{R}^{d}$, $h(s_{i})$ represents the representation of an anchor node $s_{i}$, $c(l_{j})$ represents the representation of the $j$-th class label and $\alpha$ is a hyper-parameter. The neighbourhood of a super node $p_{ij}$ is defined as the neighbourhood of its corresponding anchor node $s_{i}$, i.e., $\mathcal{N}(p_{ij})=\mathcal{N}(s_{i})$. We fuse class labels with anchor nodes, aiming to leverage the class labels to assist anchor nodes in identifying and aggregating relevant information within their neighborhoods. We then employ virtual super nodes to construct unified structural characteristics for graph data from different domains.
Encoder. We construct virtual edges $\bar{e}_{pu}$ between each super node $p$ and its neighbor nodes $u\in\mathcal{N}(p)$, and then learn weights for these edges with the attention mechanism. Let the weight $a_{pu}$ denote the importance of the edge $\bar{e}_{pu}$, i.e., the importance of neighbor node $u$ to super node $p$. The weights can be computed as follows:
$a_{pu}=\frac{exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(u)\right]))}{{\textstyle\sum_{j\in\mathcal{N}(p)}exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(j)\right])}},$ (4)
where $h(u)=x(u)$ represents the node features encoded by the LM, and $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ and $\vec{g}\in\mathbb{R}^{2d}$ are the learnable parameters. Then, we aggregate information from neighborhoods for each super node $p$ to get the corresponding representation. The final representation of super node $p_{ij}$ can be computed as follows:
$\hat{h}(p_{ij})=W_{3}\left[\beta h(p_{ij})+(1-\beta){\textstyle\sum_{u\in\mathcal{N}(p_{ij})}a_{(p_{ij})u}h(u)}\right],$ (5)
where $W_{3}\in\mathbb{R}^{d\times d}$ is the learnable parameters and $\beta$ is a hyper-parameter. Instead of relying on the original structure, we introduce super nodes and connect them to all nodes within their neighborhood by virtual edges. This new structure allows for effective information aggregation while unifying structure space across domains, thus boosting the model generalizability.
4.3 Pre-training
To optimize the pre-trained model, we leverage a contrastive loss function $\mathcal{L}_{pre}$. The super node $p$ aggregates information from its neighbor node $u\in\mathcal{N}(u)$ based on learned weights $a_{pu}$. Intuitively, neighbors that are more similar to the super node are more important. Therefore, for each encoded super node $\hat{h}(p_{ij})$, the original super node $h(p_{ij})$ before encoding is treated as a positive sample. Meanwhile, encoded super nodes from different classes are selected as negatives. We formulate the loss as the normalized temperature-scaled cross-entropy loss [1] as follows:
$\mathcal{L}_{pre}=-\frac{1}{C}\sum_{j}^{C}\frac{1}{n}\sum_{i}^{n}log\frac{exp(sim(\hat{h}(p_{ij}),h(p_{ij})/\tau)}{\sum_{q=1,q\neq j}^{C}(exp(sim(\hat{h}(p_{ij}),\hat{h}(p_{iq}))/\tau))}$ (6)
where $C$ is the number of classes, $n$ is the number of instances, $\tau$ is a temperature hyper-parameter and sim(·) is implemented with cosine similarity.
Similarity Matching & Zero-shot learning. We employ similarity matching to select the final representation for each input graph/node. Let $z_{i}$ denote the final representation of $i$-th input graph/node, it can be computed as:
$z_{i}=\hat{h}(p_{iw}),$ (7)
where $w=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j})))$, $j\in\left[1,C\right]$. Intuitively, super nodes that effectively aggregate information within their neighborhoods are most similar to their corresponding class labels. Therefore, the index $w$ based on the maximum similarity calculation is also the predicted class for node or graph classification tasks. Then, we define the prediction for the $i$-th graph/node as:
$\hat{y}_{i}=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j}))),j\in\left[1,C\right].$ (8)
Since predictions based on similarity matching do not rely on labels, and the pre-trained model can generalize to unseen data. We treat the similarity matching process as zero-shot learning. It is worth noting that when evaluating the model performance, we keep the parameters of the pre-trained model frozen.
For link prediction tasks in the zero-shot setting, we compare the similarity of input node pairs. Pairs with similarity greater than a threshold $T$ are classified as having an edge; otherwise, they are considered not to have an edge. Let $\hat{e}_{ij}$ denote the link prediction result between the pair of nodes $(v_{i},v_{j})$. It can be computed as follows:
$\hat{e}_{ij}=\begin{cases}1,\text{if}~{}sim(z_{i},z_{j})>T\\ 0,\text{otherwise}\end{cases}$ (9)
4.4 Tuning for Downstream Tasks
Next, we show how we freeze the pre-trained model and adapt it to different downstream tasks on graphs. BooG uses MLP as the classifier to predict the class of instances. Formally, the final prediction for $i$-th input graph/node is given by:
$\hat{y_{i}}=\arg\max(MLP(z_{i})).$ (10)
In link prediction tasks, we standardize it as a binary classification task and employ the concatenated representation of node pairs as the input to the MLP for prediction. In both supervised and few-shot learning scenarios, we employ cross-entropy loss to train the MLP for all tasks.
4.5 Time Complexity
We analyze the time complexity of BooG in this section. For BooG, we denote the maximum sequence length of node textual feature as $L$ and the number of instances as $n$. For an input graph/node, the complexity is $O(L^{2}d+Ld^{2})$, where $d$ is the dimensionality of the embeddings. The time complexity is $O(n*(L^{2}d+Ld^{2}))$ for processing $n$ instances. The encoder’s time complexity primarily depends on the feature dimensionality and the structure of the graph. The time complexity of the attention mechanism can be expressed as $O(2d^{2}n+2dn)$. Next, the complexity of message aggregation is $O(\bar{k}dn+d^{2}n)$, where $\bar{k}$ is the average number of neighbors per node. The overall time complexity of encoder is $O(k_{1}d^{2}n+k_{2}dn)$, where $k_{1},k_{2}$ are the coefficients. For similarity matching, the time complexity is $O(dnC)$, where $C$ is the number of classes. Overall, our method BooG has a linear time complexity w.r.t. $n$.
5 Experiments
In this section, we evaluate the generalizability of BooG on both node-level and graph-level tasks.
5.1 Experimental Settings
Datasets. We conduct experiments on six benchmark datasets, which include 4 node-level graph datasets (Cora, Pubmed, ogbn-arxiv, Wiki-CS) and 2 graph-level datasets (PCBA, HIV). Due to the space limitation, we move details on datasets (Sec. A), experimental setup (Sec. B) and hyper-parameter sensitivity analysis (Sec. C) to Appendix.
Baselines. We compare BooG with six baselines, which can be categorized into four types. (1) MLP utilizes a multi-layer perceptron to extract deep features individually for each node. (2) Graph Neural Networks: GAT [25] computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN [29] enhances the representation power of GNNs by employing a distinct graph encoding method that emphasizes the discrimination of non-isomorphic structures. (3) Graph Pre-training Models: GraphCL [30] utilizes pre-training of graph models through the application of a self-discriminative contrastive learning task on learned node embeddings. DGI [26] introduces a self-supervised pre-training task that aims to maximize the mutual information between the local node view and the global graph view. (4) Graph foundation models: OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs.
5.2 Performance Results
We next evaluate the model generalizability across various learning scenarios and downstream tasks. We employ TAGs for all methods to unify all input data. Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models. BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora. In the supervised learning setting, the parameters of pre-trained models are frozen and the MLP is trained in a supervised manner. Other baselines are trained directly on the target datasets. In the few-shot learning setting, we freeze the parameters of the pre-trained methods and train their MLPs using a $N$-way $K$-shot task. For other methods, we train them using the $N$-way $K$-shot task, i.e., they are directly trained on target datasets with labels. To accomplish this, we construct a support set by randomly selecting $K$ examples per class from the training split. For zero-shot learning, different from the above two scenarios, OFA is first pre-trained on the citation network Cora. For GraphCL and DGI, predictions are generated by comparing the similarity between the instance representations derived from their models and the class label representations encoded by the LM. Finally, we evaluate all methods on the test sets of the target datasets.
Performance analysis. Table 1, Table 2 and Table 3 report the results in supervised learning, few-shot learning, and zero-shot learning, respectively. From these tables, we observe that: (1) BooG demonstrates the best results in most cases. (2) The GNN models GAT and GIN perform better in specific domains in supervised learning and few-shot learning. For example, GAT excels on citation network datasets, while GIN is more suited to molecular graphs. However, these methods perform poorly in other domains because they are typically designed for a single graph domain, especially in few-shot learning. (3) The two self-supervised learning methods, GraphCL and DGI, exhibit poor domain generalization capabilities in few-shot learning and zero-shot learning, since they only use the raw graph structure. (4) OFA demonstrates a certain degree of domain generalization capabilities in few-shot learning, highlighting the effectiveness of in-context learning. However, BooG outperforms OFA in most cases because it constructs unified structural characteristics for graphs from different domains, thereby enhancing the model generalizability. Overall, BooG shows strong generalization ability in all scenarios, achieving results comparable to or even better than baseline methods, which proves its effectiveness.
Table 1: Experiment results in supervised learning. We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks. We highlight the best score on each dataset in bold and the runner-up score with underline.
Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
[…]
Table 2: Experiment results in few-shot learning. We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks. We highlight the best score on each dataset in bold and the runner-up score with underline.
Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
[…]
Table 3: Experiment results in zero-shot learning. We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks. We highlight the best score on each dataset in bold and the runner-up score with underline. Since OFA is first pre-trained on the citation network Cora with training labels, so we did not report the results of OFA on Cora in zero-shot learning.
Datasets and tasks Cora Pubmed ogbn-arxiv Wiki-CS Cora-link PCBA HIV
[…]
DGI 38.36 28.46 24.11 23.76 35.85 21.02 25.11
[…]
Figure 4: t-SNE visualization on Cora.
Figure 5: Ablation study.
5.3 Visualization
We further evaluate the expressive power of BooG. We visualize node representations before and after pre-training on Cora with the t-SNE technique [24]. Colors are used to distinguish different node classes. Before pre-training, the node representations are encoded by the LM. As shown in Figure 5, the node representations of different classes are more clearly separated after pre-training. This indicates that BooG can learn more expressive representations.
5.4 Ablation Study
The ablation study is done to understand the importance of main components of BooG in supervised learning and 5-shot learning scenarios on Cora and HIV. The model is pre-trained on Cora. We first remove the class labels from the super nodes and call this variant BooG-nc (no class label). Accordingly, in the pre-training task, we use the anchor node before aggregation as the positive sample and treat other instances of dataset as negative samples. Secondly, we remove the anchor nodes from the super nodes and treat the anchor node as a member of the contextual neighborhood. We call this variant BooG-na (no anchor node). Further, we remove the similarity matching process and concatenate the representations based on different class labels as the final representation of the input to the MLP. We call this variant BooG-ns (no similarity matching). We compare BooG with these three variants, and the results are presented in Figure 5. Our findings show that BooG outperforms all the variants on the two datasets. Further, the decline in performance exhibited by both BooG-nc and BooG-na highlights the importance of super nodes. In other words, the super nodes significantly enhance model generalizability, especially in graph domain transfer.
6 Conclusion
In this paper, we boost graph foundation model from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified. Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges. This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains. In this way, cross-domain model generalization can be boosted from structural perspective. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Our extensive experiments and ablation studies validate the effectiveness of BooG.
Appendix A Datasets
In this section, we introduce the datasets used in the experiments. The statistics of datasets are listed in Table 4, and we describe each dataset in detail below:
Cora is a homophilous graphs commonly used as benchmarks. It consists of 2708 scientific publications classified into one of 7 classes. In this dataset, each node represents a scientific paper and each edge represents a citation. We collect raw text from [5].
Pubmed is a homophilous graphs commonly used as benchmarks. It consists of 19,717 scientific publications from PubMed database pertaining to diabetes classified into one of 3 classes. We collect raw text from [5].
ogbn-arXiv is an academic citation network graph dataset, consisting of papers from arXiv and their citation relationships. Each node represents a paper, and edges represent the citation relationships between papers. The task is to predict the 40 subject areas of arXiv CS papers. We collect raw text from [8].
Wiki-CS is an Internet link network where each node represents a Wikipedia page and each edge represents a reference link. Each node’s label corresponds to the category of the entry. We collect raw text from [14].
PCBA is a large-scale chemical bioactivity dataset used for graph machine learning research in cheminformatics and drug discovery. Each node represents an atom of a molecule, and the edges represent the chemical bonds between atoms within the molecule. The primary tasks in the PCBA dataset are to predict the activity of chemical molecules across various bioactivity assays. We collect raw text from [14].
HIV is a subset of the PubChem BioAssay dataset, focusing on predicting the activity of chemical molecules against the HIV virus. The task in the HIV dataset involves predicting the activity of chemical molecules against the HIV virus. We collect raw text from [14].
For Cora and PubMed, we measure the performance of all models on the test sets over 10 random splits as suggested in [20] and report the average accuracy. For ogbn-arxiv, we follow the official splits [8]. Following the experimental procedure suggested by OGB, we repeat each experiment for 10 times with random seeds and report the average accuracy. For Wiki-CS, we follow the official splits [19] with 20 different training splits and report the average accuracy. For PCBA and HIV, we follow the official splits [32], repeat each experiment for 10 times with random seeds and report the average accuracy. The statistics and details of these datasets can be found in Table 4.
Table 4: Statistics of all text-attributed graph datasets.
Dataset Domain Task #Graphs Avg.#Nodes Avg.#Edges Raw Texts
Cora Citation Node/Link 1 2,708 10,556 paper titles and abstracts
Pubmed Citation Node/Link 1 19,717 44,338 paper titles and abstracts
ogbn-arxiv Citation Node 1 169,343 1,166,243 paper titles and abstracts
Wiki-CS Web link Node 1 11,701 216,123 wikipedia entry names and contents
PCBA Molecule Graph 437,929 26.0 28.1 textual descriptions of atoms/bonds
HIV Molecule Graph 41,127 25.5 27.5 textual descriptions of atoms/bonds
Appendix B Experimental Setup
We implement BooG by PyTorch and conducted the experiments with one A100 GPU. The model is optimized by Adam [10]. We employ TAGs for all methods to unify all input data. Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models. BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora. For other self-supervised learning baseline methods, we train a linear classifier on the top of the embeddings derived from the frozen model. Then, we evaluate the model performance on the test sets of all six datasets with different tasks. We perform a grid search to tune hyper-parameters based on the validation set. Details on the search space of these hyper-parameters can be found in Table 5.
Table 5: Grid search space.
Hyper-parameter Search space
lr {0.01, 0.02, 0.1, 0.2}
dropout [0.0,0.9]
weight decay {5e-5, 1e-5, 5e-4, 1e-4, 5e-3}
$\alpha$ [0.1,0.9]
$\beta$ [0.1,0.9]
$\tau$0.1,1,10}
T [0.1,0.9]
Appendix C Sensitivity Analysis
Here we study the sensitivity of two key hyper-parameters: $\alpha$ and $\beta$. The results are shown in Figure 6. From the figure, we see that the hyper-parameter $\alpha$, which represents the importance of class labels when constructing super nodes, has a significant impact on the model’s performance. If the weight of the class labels is too small, the class labels will not be effective, leading to a decline in model performance. Conversely, if the weight of the class labels is too large, the representations of the instances may lose their individuality, also resulting in a performance drop. For hyper-parameter $\beta$, BooG consistently performs well across a wide range of values, which demonstrates that $\beta$ does not significantly impact the performance of BooG.
Figure 6: Hyper-parameter sensitivity analysis.
Appendix D Broader Impacts
The primary objective of this study is to chieves robust performance and generalization capabilities across various graph domains and tasks with a graph foundation model. The results indicate that the proposed model contribute to enhance the model generalizability, which can have significant implications in various fields such as social network analysis and molecular property prediction. However, this work may also have some negative impacts. As deep neural network models become more widespread, they can be prone to misuse, potentially resulting in detrimental outcomes. Additionally, the privacy of training data remains an issue, as sensitive information may be unintentionally leaked through the model. We could incorporate responsible AI techniques for GNNs into our model to avoid the issues of misuse and privacy.
Appendix E Limitations
We discuss the limitations and future directions of our work in this section. This method is currently only suitable for graph classification, node classification and link prediction tasks. In future work, we aim to extend our method to a wider range of graph tasks. In addition, our method does not take advantage of the textual attributes of the edges, and we also aim to extend our method to more data attributes.